# Cross-modal Full-mode Fine-grained Alignment for Text-to-Image Person Retrieval

### v1

@@ Abstract.
Text-to-Image Person Retrieval (TIPR) is a cross-modal matching task that aims to retrieve the most relevant person images based on a given text query. The key challenge in TIPR lies in achieving effective alignment between textual and visual modalities within a common latent space. To address this challenge, prior approaches incorporate attention mechanisms for implicit cross-modal local alignment. However, they lack the ability to verify whether all local features are correctly aligned. Moreover, existing methods primarily focus on hard negative samples during model updates, with the goal of refining distinctions between positive and negative pairs, often neglecting incorrectly matched positive pairs. To alleviate these issues, we propose FMFA, a cross-modal Full-Mode Fine-grained Alignment framework, which enhances global matching through explicit fine-grained alignment and existing implicit relational reasoning—hence the term “full-mode”—without requiring additional supervision. Specifically, we design an Adaptive Similarity Distribution Matching (A-SDM) module to rectify unmatched positive sample pairs. A-SDM adaptively pulls the unmatched positive pairs closer in the joint embedding space, thereby achieving more precise global alignment. Additionally, we introduce an Explicit Fine-grained Alignment (EFA) module, which makes up for the lack of verification capability of implicit relational reasoning. EFA strengthens explicit cross-modal fine-grained interactions by sparsifying the similarity matrix and employs a hard coding method for local alignment. Our proposed method is evaluated on three public datasets, achieving state-of-the-art performance among all global matching methods. Our code is available at https://github.com/yinhao1102/FMFA.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: tommjournalvolume: 1journalnumber: 1article: 1publicationmonth: 1ccs: Information systems Image searchccs: Computing methodologies Object identification
@@ 1. Introduction
Text-to-Image Person Retrieval (TIPR) aims to comprehend natural language descriptions and retrieve the most relevant person image from a large image gallery (Lu et al., 2019). Unlike general image-text retrieval (Sogi et al., 2024; Chen et al., 2024, 2023), which tends to achieve semantic-based matching between text and image, TIPR is specifically designed for identifying individuals. TIPR requires the accurate modeling of fine-grained correspondences between textual and visual modalities, owing to the large intra-class variance and small inter-class difference. This substantial intra-class variation arises from two aspects: (1) visual appearances of the same identity exhibit dramatic variations under different poses, viewpoints, and illumination conditions, and (2) textual descriptions are influenced by differences in phrasing, word order and textual ambiguities. Therefore, the primary challenges in TIPR are how to extract discriminative global representations from image-text pairs and how to achieve precise cross-modal fine-grained alignment. Existing methods for tackling these challenges can be roughly divided into two main categories: global matching methods and local matching methods.
@@
Some global matching methods (Zhang and Lu, 2018; Zheng et al., 2020) obtain discriminative global representations by aligning images and texts, which are projected into a joint embedding space. Their commonly used loss functions include the Cross-Modal Projection Matching (CMPM) loss (Zhang and Lu, 2018) and the Similarity Distribution Matching (SDM) loss (Jiang and Ye, 2023). The CMPM loss focuses on the difference between the scalar projection of image-text pairs and their corresponding matching labels. In comparison, the SDM loss enhances global image-text matching by minimizing the Kullback-Leibler (KL) divergence between the normalized similarity score distribution of image-text pairs and the ground-truth label distribution. In addition, the SDM loss incorporates a temperature hyperparameter to make model updates concentrate on hard negative samples, yet it leads to the neglect of unmatched positive pairs, as shown in Figure 1(a). However, in TIPR, the accurate matching of positive pairs is prioritized over merely distinguishing between positive and negative pairs. Meanwhile, some local matching methods (Bai et al., 2023b; Park et al., 2024; Ergasti et al., 2025) incorporate attention mechanisms to achieve cross-modal fine-grained alignment. For instance, RaSa (Bai et al., 2023b) constructs a cross-modal encoder to generate multimodal representations for subsequent fine-grained alignment. Building on RaSa, MARS (Ergasti et al., 2025) integrates an MAE decoder (He et al., 2022) to reconstruct masked image patch sequences into their original unmasked form, thereby facilitating cross-modal fine-grained alignment. However, these methods rely on attention mechanisms to implicitly aggregate local image-text representations. As a result, they yield only the final multimodal representation, without revealing the details of the aggregation process. Consequently, these implicit aggregation methods make it difficult to determine whether the aggregated multimodal representations correctly encode the corresponding visual and textual information.
@@
To overcome these issues, we introduce FMFA, a cross-modal Full-Mode Fine-grained Alignment framework, which enhances global matching through full-mode fine-grained alignment, including explicit fine-grained image-text alignment and existing implicit relational reasoning. Specifically, we design an Adaptive Similarity Distribution Matching (A-SDM) module to ensure the correct matching of positive image-text pairs. Within the joint embedding space, the A-SDM module adaptively pulls positive pairs closer together. In cases of mismatched positive pairs, the A-SDM module adaptively regulates the pulling force based on their relative distance within the joint embedding space, as shown in Figure 1(b), thus improving cross-modal global alignment. Inspired by the observation that multiple image patches usually correspond to one word in the caption (Bica et al., 2024), we introduce an Explicit Fine-Grained Alignment (EFA) module. The EFA module derives multimodal representations through explicit aggregation with a sparse similarity matrix. During this process, the sparse similarity matrix between text and image reflects the contribution of textual and visual representations to the final multimodal representation. To minimize redundancy and reduce the computational cost during training, the EFA module employs hard coding alignment between the aggregated multimodal representation and its original visual and textual representations. These designs enable EFA to achieve cross-modal fine-grained interactions and help the backbone network to extract more discriminative global image-text representations without requiring additional supervision. FMFA is evaluated on three public benchmarks (Li et al., 2017; Ding et al., 2021; Zhu et al., 2021), and it achieves state-of-the-art performance while maintaining high inference speed. Our main contributions are summarized as follows:
  • We propose FMFA to explicitly exploit fine-grained interactions to enhance cross-modal alignment without additional supervision and inference cost.
  • We introduce an adaptive similarity distribution matching module, which is designed to accurately align image-text pairs within a joint embedding space. It adaptively adjusts to narrow the distance between mismatched positive pairs, ensuring more precise matching.
@@ 2. Related Work
Text-to-Image Person Retrieval (TIPR) was first introduced by Li et al. (Li et al., 2017), who constructed the CUHK-PEDES dataset. The core challenge of TIPR is to learn a common latent space where visual and textual representations are well aligned. Existing methods can be broadly categorized into global and local matching approaches.
@@
Early global methods (Shu et al., 2022; Zheng et al., 2020; Zhu et al., 2021) directly aligned the global representations of images and text in a joint embedding space. Schroff et al. (Schroff et al., 2015) proposed a triplet ranking loss to enforce a margin constraint between positive and negative pairs, and Zhang et al. (Zhang and Lu, 2018) introduced the CMPM/C loss to minimize the discrepancy between the scalar projection of image-text pairs and their labels. However, these global methods lack cross-modal fine-grained interactions, which restrict their ability to capture detailed semantic correspondences. To address this limitation, early local matching methods (Wang et al., 2020; Gao et al., 2021; Shao et al., 2022) explicitly aligned local visual and textual features to achieve fine-grained cross-modal interactions. Nevertheless, they rely on unimodal pre-trained models (e.g., BERT (Devlin et al., 2019) and ResNet (He et al., 2016)), failing to exploit the strong cross-modal alignment capability of recent pre-trained Vision-Language Models (VLMs) (Li et al., 2022; Yan et al., 2023a; Li et al., 2021).
@@
Recent local matching methods (Ergasti et al., 2025; Lu et al., 2025; Qin et al., 2024; Wu et al., 2024; Huang et al., 2025) have benefited greatly from VLMs and introduced VLMs to enhance cross-modal alignment. Park et al. (Park et al., 2024) utilized a modified Contrastive Language-Image Pre-training (CLIP) (Radford et al., 2021) model as the feature extractor and designed a slot attention-based (Locatello et al., 2020) part discovery module to identify discriminative human parts without extra supervision, while Bai et al. (Bai et al., 2023b) used the align-before-fuse model (Li et al., 2021) as the backbone and introduced a cross-modal encoder for fine-grained alignment. Although effective, these methods involve complex computations during inference, leading to high time and memory costs, which limit their applicability to real-time systems.
@@
On another line of research, several studies (Yang et al., 2023; Shao et al., 2023; Tan et al., 2024) have explored leveraging large-scale image-text pairs in the person Re-IDentification (ReID) domain to VLMs. Zuo et al. (Zuo et al., 2024) utilized CUHK-PEDES and ICFG-PEDES to train an image captioner, aiming to generate comprehensive textual descriptions for pedestrian images. Yang et al. (Yang et al., 2023) employed BLIP-2 (Li et al., 2023) to produce attribute-aware captions for diffusion-generated pedestrian images (Rombach et al., 2022), while Jiang et al. (Jiang et al., 2025) leveraged recent multi-modal large language models (MLLMs), such as Qwen-VL (Bai et al., 2023a) and LLaVA (Liu et al., 2024), to automatically annotate large-scale ReID datasets in a human-like manner. The CLIP models pre-trained on large-scale ReID datasets exhibit strong zero-shot performance. Their compatibility with global matching methods—which relies solely on global features and has a simple inference pipeline—makes them particularly suitable for direct fine-tuning in such settings.
Among the global matching methods, IRRA (Jiang and Ye, 2023) recently attempted to enhance fine-grained interactions by introducing an Implicit Relation Reasoning (IRR) module, utilizing attention mechanisms to capture cross-modal relations. However, this implicit alignment cannot guarantee that all corresponding local features are correctly matched. In light of these limitations, we propose FMFA, which aims to enhance the global matching ability of the model by achieving cross-modal full-mode fine-grained alignment, including explicit fine-grained alignment and implicit relation reasoning.
@@
Figure 2. Overview of the proposed FMFA framework. (a) Our FMFA consists of a dual-stream feature extraction backbone and four representation learning branches, i.e. Identity Identification (ID loss), Adaptive Similarity Distribution Matching (A-SDM), Explicit Fine-grained Alignment (EFA) and Implicit Relation Reasoning (IRR). The former two are coarse-grained alignment modules and the latter two are fine-grained alignment modules. Modules connected by dashed lines will be removed during inference stage. (b) A-SDM obtains the adaptive weight through the similarity score of unmatched positive pairs, adaptively pulling positive pairs closer and dynamically adjusting the pulling force. (c) EFA achieves fine-grained interaction by hard coding alignment of token or patch embeddings and their corresponding language-grouped vision embedding, which is derived from an aggregated sparse similarity matrix.
@@ 3. Method
In this section, we present our proposed FMFA framework. The overview of FMFA is illustrated in Figure 2 and the details are introduced in the following subsections.
@@ 3.1. Feature Extraction
Inspired by the success of IRRA (Jiang and Ye, 2023), we use the modified full CLIP (Radford et al., 2021) visual and textual encoders to enhance cross-modal alignment capabilities while reducing inference costs.
Visual Modality. Given an input image $I\in R^{H\times W\times C}$, a CLIP pre-trained ViT model is adopted to obtain the image embeddings. We first split $I$ into a sequence of $N=H\times W/P^{2}$ fixed-sized non-overlapping patches, where $P$ denotes the patch size, and then map the patch sequences to 1D tokens $\left\{f_{i}^{v}\right\}|_{i=1}^{N}$ by a trainable linear projection. With the injection of the positional embedding and the extra [CLS] token, the sequence of tokens $\left\{f_{cls}^{v},f_{1}^{v},...,f_{N}^{v}\right\}$ are input into $L$-layer transformer blocks to model the correlations of each patch. Finally, a linear projection is adopted to map $f_{cls}^{v}$ to the joint image-text embedding space, which serves as the global image representation.
Textual Modality. For an input text $T$, we directly use the CLIP-Xformer textual encoder (Radford et al., 2021) to extract the text embeddings. Following CLIP, the lower-cased byte pair encoding (Sennrich et al., 2016) is firstly employed to tokenize the input text description. The text description is enclosed with [SOS] and [EOS] tokens to indicate the start and end of the sequence. The tokenized text $\{f_{sos}^{t},f_{1}^{t},...,f_{eos}^{t}\}$ is then fed into the textual encoder and exploits the correlations of each patch by masked self-attention. The output of the highest layer of the transformer at the [EOS] token, $f_{eos}^{t}$, is linearly projected into the image-text joint embedding space to obtain the global text representation.
@@ 3.2. Adaptive Similarity Distribution Matching
Adopted from IRRA (Jiang and Ye, 2023), we introduce a novel Adaptive Similarity Distribution Matching (A-SDM) module, which aims to adaptively pull the unmatched positive image-text pairs into a joint embedding space, thereby enhancing the cross-modal global matching capability of model.
Given a mini-batch of $B$ image-text pairs, for each global text representation $g_{i}^{t}$ and global image representation $g_{j}^{v}$, we construct a set of image-text representation pairs as $\{(g_{i}^{t},g_{j}^{v}),y_{i,j}\}_{j=1}^{N}$, where $y_{i,j}$ is the matching label. $y_{i,j}=1$ means that $(g_{i}^{t},g_{j}^{v})$ is a matched pair, while $y_{i,j}=0$ means that $(g_{i}^{t},g_{j}^{v})$ is an unmatched pair. Let $cos(\mathbf{u},\mathbf{v})=\mathbf{u}^{\top}\mathbf{v}/\|\mathbf{u}\|\|\mathbf{v}\|$ denotes the cosine similarity of $u$ and $v$. Then, the similarity matrix of matching pairs can be simply calculated by the following softmax function:
(1) $p_{i,j}=\frac{exp(cos(g_{i}^{t},g_{j}^{v})/\tau_{1})}{\sum_{k=1}^{N}exp(cos(g_{i}^{t},g_{k}^{v})/\tau_{1})},$
where $\tau_{1}$ is a temperature hyperparameter which controls the probability distribution peaks. The matching probability $p_{i,j}$ can be viewed as the proportion of the cosine similarity score between $g_{i}^{t}$ and $g_{j}^{v}$ to the sum of cosine similarity scores between $g_{i}^{t}$ and $\{g_{j}^{v}\}_{j=1}^{N}$ in a mini-batch.
Let the ithi-th text $T_{i}$ in the mini-batch be the query text and $I_{i}$ be the corresponding image for $T_{i}$ at rank-$k$, where $k>1$. Different from IRRA (Jiang and Ye, 2023), we propose to calculate adaptive weight based on the similarity between the query text $T_{i}$ and all image representations:
@@
where $\alpha$ is a weight factor which reflects the contribution of unmatched image-text pairs to the cross-modal global matching capability of the model. $w_{i}^{t2i}=1$ indicates that the query text $T_{i}$ and its corresponding image $I_{i}$ are successfully matched, while $w_{i}^{t2i}>1$ indicates that $T_{i}$ and $I_{i}$ are unmatched. The A-SDM loss from text to image within a mini-batch is then computed as:
@@
where $\epsilon$ is a small constant to avoid numerical problems, and $q_{i,j}=y_{i,j}/\sum_{k=1}^{B}$ is the true matching probability.
Symmetrically, the A-SDM loss from image to text $\mathcal{L}_{i2t}$ can be obtained by exchanging the visual and textual modalities. The bi-directional A-SDM loss is calculated by:

### v2

@@ Abstract.
Text-to-Image Person Retrieval (TIPR) is a cross-modal matching task designed to identify the person images that best correspond to a given textual description. The key difficulty in TIPR is to realize robust correspondence between the textual and visual modalities within a unified latent representation space. To address this challenge, prior approaches incorporate attention mechanisms for implicit cross-modal local alignment. However, they lack the ability to verify whether all local features are correctly aligned. Moreover, existing methods tend to emphasize the utilization of hard negative samples during model optimization to strengthen discrimination between positive and negative pairs, often neglecting incorrectly matched positive pairs. To mitigate these problems, we propose FMFA, a cross-modal Full-Mode Fine-grained Alignment framework, which enhances global matching through explicit fine-grained alignment and existing implicit relational reasoning—hence the term “full-mode”—without introducing extra supervisory signals. In particular, we propose an Adaptive Similarity Distribution Matching (A-SDM) module to rectify unmatched positive sample pairs. A-SDM adaptively pulls the unmatched positive pairs closer in the joint embedding space, thereby achieving more precise global alignment. Additionally, we introduce an Explicit Fine-grained Alignment (EFA) module, which makes up for the lack of verification capability of implicit relational reasoning. EFA strengthens explicit cross-modal fine-grained interactions by sparsifying the similarity matrix and employs a hard coding method for local alignment. We evaluate our method on three public datasets, where it attains state-of-the-art results among all global matching methods. The code for our method is publicly accessible at https://github.com/yinhao1102/FMFA.
@@
copyright: acmlicensedjournal: TOMMjournalyear: 2025journalvolume: 1journalnumber: 1article: 1publicationmonth: 1doi: 10.1145/3786798ccs: Information systems Image searchccs: Computing methodologies Object identification
@@ 1. Introduction
Text-to-Image Person Retrieval (TIPR) seeks to understand natural language descriptions and identify the most relevant person image within a large gallery (lu2019vilbert). Unlike general image-text retrieval (wang2017adversarial; sogi2024object; chen2024make; chen2023vilem; wang2025geometric), which tends to achieve semantic-based matching between text and image, TIPR is specifically designed for identifying individuals. TIPR requires the accurate modeling of fine-grained correspondences between textual and visual modalities, owing to the large intra-class variance and small inter-class difference. This substantial intra-class variation arises from two aspects: (1) visual appearances of the same identity exhibit dramatic variations under different poses, viewpoints, and illumination conditions, and (2) textual descriptions are influenced by differences in phrasing, word order and textual ambiguities. Therefore, the primary challenges in TIPR are how to extract discriminative global representations from image-text pairs and how to achieve precise cross-modal fine-grained alignment. Existing methods for tackling these challenges can be roughly divided into two main categories: global matching methods and local matching methods.
@@
Some global matching methods (zhang2018deep; zheng2020dual) obtain discriminative global representations by aligning images and texts, which are projected into a joint embedding space. Their widely adopted loss functions include the Cross-Modal Projection Matching (CMPM) loss (zhang2018deep) and the Similarity Distribution Matching (SDM) loss (jiang2023cross). The CMPM loss highlights the gap between the scalar projections of image-text pairs and their matched label indicators. In comparison, the SDM loss boosts global matching performance by minimizing the Kullback-Leibler (KL) divergence between the normalized similarity profile of image-text pairs and the true label distribution. In addition, the SDM loss incorporates a temperature hyperparameter to make model updates concentrate on hard negative samples, yet it leads to the neglect of unmatched positive pairs, as shown in Figure 1(a). However, in TIPR, the accurate matching of positive pairs is prioritized over merely distinguishing between positive and negative pairs. Meanwhile, some local matching methods (bai2023rasa; park2024plot; ergasti2025mars) incorporate attention mechanisms to achieve cross-modal fine-grained alignment. For instance, RaSa (bai2023rasa) constructs a cross-modal encoder to generate multimodal representations for subsequent fine-grained alignment. Building on RaSa, MARS (ergasti2025mars) integrates a Masked AutoEncoder (MAE) decoder (he2022masked) to reconstruct masked image patch sequences into their original unmasked form, thereby facilitating cross-modal fine-grained alignment. However, these methods rely on attention mechanisms to implicitly aggregate local image-text representations. As a result, they yield only the final multimodal representation, without revealing the details of the aggregation process. Consequently, these implicit aggregation methods make it difficult to determine whether the aggregated multimodal representations correctly encode the corresponding visual and textual information.
@@
To remedy these concerns, we propose FMFA, a cross-modal Full-Mode Fine-grained Alignment framework, which enhances global matching through full-mode fine-grained alignment, including explicit fine-grained image-text alignment and existing implicit relational reasoning. Specifically, we design an Adaptive Similarity Distribution Matching (A-SDM) module to ensure the correct matching of positive image-text pairs. Within the joint embedding space, the A-SDM module adaptively pulls positive pairs closer together. In cases of mismatched positive pairs, the A-SDM module adaptively regulates the pulling force based on their relative distance within the joint embedding space, as shown in Figure 1(b), thus improving cross-modal global alignment. Based on the insight that each word in a caption can be associated with several image patches (bica2024improving), we introduce an Explicit Fine-Grained Alignment (EFA) module. The EFA module derives multimodal representations through explicit aggregation with a sparse similarity matrix. During this process, the sparse similarity matrix between text and image reflects the contribution of textual and visual representations to the final multimodal representation. To minimize redundancy and reduce the computational cost during training, the EFA module employs hard coding alignment between the aggregated multimodal representation and its original visual and textual representations. These designs allow EFA to realize fine-grained cross-modal interactions and assist the backbone network in learning more distinctive global image-text representations without introducing additional supervision. FMFA is evaluated on three public benchmarks (li2017person; ding2021semantically; zhu2021dssl), and it attains competitive top-level performance along with high inference efficiency. We highlight our key contributions below:
  • We introduce FMFA to explicitly leverage fine-grained interactions for improving cross-modal alignment, without incurring extra supervision or inference overhead.
  • We present an adaptive similarity distribution matching module aimed at precisely aligning image-text pairs in a shared embedding space. It adaptively adjusts to narrow the distance between mismatched positive pairs, ensuring more precise matching.
@@ 2. Related Work
Text-to-Image Person Retrieval (TIPR) was initially proposed by Li et al. (li2017person), who created the CUHK-PEDES dataset. Unlike visual-based person retrieval (cheng2022hybrid; cheng2025semantic; he2025exploring; yang2025feature), the core challenge of TIPR lies in constructing a shared latent space that enables coherent alignment between visual and textual representations. Existing methods can be typically classified into global and local matching approaches.
@@
Early global methods (zheng2020dual; zhu2021dssl) directly aligned the global representations of images and text in a joint embedding space. Schroff et al. (schroff2015facenet) proposed a triplet ranking loss to enforce a margin constraint between positive and negative pairs, and Zhang et al. (zhang2018deep) introduced the CMPM/C loss to minimize the discrepancy between the scalar projection of image-text pairs and their labels. However, these global methods lack cross-modal fine-grained interactions, which restrict their ability to capture detailed semantic correspondences. To address this limitation, early local matching methods (wang2020vitaa; gao2021contextual; shao2022learning) explicitly aligned local visual and textual features to achieve fine-grained cross-modal interactions. Nevertheless, they rely on unimodal pre-trained models (e.g., BERT (devlin2019bert) and ResNet (he2016deep)), failing to exploit the strong cross-modal alignment capability of recent pre-trained Vision-Language Models (VLMs) (li22blip; yan2023clip; li2021align).
@@
Recent local matching methods (ergasti2025mars; lu2025prompt; qin2024noisy; wu2024text; huang2025cross) have benefited greatly from VLMs and introduced VLMs to enhance cross-modal alignment. Park et al. (park2024plot) utilized a modified Contrastive Language-Image Pre-training (CLIP) (radford2021learning) model as the feature extractor and designed a slot attention-based (locatello2020object) part discovery module to identify discriminative human parts without extra supervision, while Bai et al. (bai2023rasa) used the align-before-fuse model (li2021align) as the backbone and introduced a cross-modal encoder for fine-grained alignment. Although effective, these methods involve complex computations during inference, leading to high time and memory costs, which limit their applicability to real-time systems.
@@
On another line of research, several studies (yang2023towards; shao2023unified; tan2024harnessing) have explored leveraging large-scale image-text pairs in the person Re-IDentification (ReID) domain to VLMs. Zuo et al. (zuo2024plip) utilized CUHK-PEDES and ICFG-PEDES to train an image captioner, aiming to generate comprehensive textual descriptions for pedestrian images. Yang et al. (yang2023towards) employed BLIP-2 (li2023blip2) to produce attribute-aware captions for diffusion-generated pedestrian images (rombach2022high), while Jiang et al. (jiang2025modeling) leveraged recent Multi-modal Large Language Models (MLLMs), such as Qwen-VL (bai2023qwenvl) and LLaVA (liu2024llavanext), to automatically annotate large-scale ReID datasets in a human-like manner. The CLIP models pre-trained on large-scale ReID datasets exhibit strong zero-shot performance. Their compatibility with global matching methods—which relies solely on global features and has a simple inference pipeline—makes them particularly suitable for direct fine-tuning in such settings.
Recent global matching methods (shu2022see; he23vgsg; jiang2023cross) have integrated local fine-grained alignment modules into global matching frameworks to obtain more discriminative global representations. Shu et al. (shu2022see) introduced a bidirectional mask modeling mechanism that randomly masks image patches and text words, encouraging the model to infer missing semantics and implicitly learn local visual-textual correspondences. He et al. (he23vgsg) proposed the Vision-Guided Semantic-Group (VGSG) network to cluster textual tokens into semantic groups and align them with corresponding visual regions under the guidance of vision features, achieving group-level fine-grained alignment within a global representation space. Similarly, Jiang et al. (jiang2023cross) developed IRRA to employ an Implicit Relation Reasoning (IRR) module based on attention mechanisms to capture latent cross-modal relations, enhancing global alignment. Although these methods enhance fine-grained cross-modal interactions within global matching frameworks, their implicit or group-level alignment strategies may still fail to guarantee precise local correspondences. In light of these limitations, we propose FMFA, which aims to enhance the global matching ability of the model by achieving cross-modal full-mode fine-grained alignment, including explicit fine-grained alignment and implicit relation reasoning.
@@
Figure 2. The illustration of our FMFA framework. (a) Our FMFA contains a two-stream feature extraction network and four distinct modules for structured representation learning, namely Identity Identification (ID loss), Adaptive Similarity Distribution Matching (A-SDM), Explicit Fine-grained Alignment (EFA) and Implicit Relation Reasoning (IRR). The former two are coarse-grained alignment modules and the latter two are fine-grained alignment modules. Modules linked via dashed connections are omitted during inference. (b) A-SDM obtains the adaptive weight through the similarity score of unmatched positive pairs, adaptively pulling positive pairs closer and dynamically adjusting the pulling force. (c) EFA achieves fine-grained interaction by hard coding alignment of token or patch embeddings and their corresponding language-grouped vision embedding, which is derived from an aggregated sparse similarity matrix.
@@ 3. Method
This section introduces the proposed FMFA framework. Figure 2 presents an overview of FMFA, and further details of the framework are elaborated in the subsequent subsections.
@@ 3.1. Feature Extraction
Motivated by the success of IRRA (jiang2023cross), we use the modified full CLIP (radford2021learning) visual and textual encoders to enhance cross-modal alignment capabilities while reducing inference costs.
Visual Modality. Given an input image $I\in\mathbb{R}^{H\times W\times C}$, we employ a CLIP-pretrained Vision Transformer (ViT) to attain its image representation. An image is first divided into $N=H\times W/P^{2}$ distinct patches of size $P\times P$, which are then transformed into one-dimensional token embeddings $\{f_{i}^{v}\}_{i=1}^{N}$ via a learnable linear projection. After adding positional encodings and a [CLS] token, the sequence $\{f_{cls}^{v},f_{1}^{v},\dots,f_{N}^{v}\}$ is passed through $L$ transformer layers to capture dependencies among patches. Finally, the [CLS] token embedding $f_{cls}^{v}$ is linearly mapped into the joint image-text embedding space, producing the compact global feature of the image.
Textual Modality. Given an input text $T$, we utilize the CLIP-Xformer textual extractor (radford2021learning) to obtain its embedding. The text is first tokenized through lower-cased Byte Pair Encoding (BPE) (sennrich2015neural) and framed with [SOS] and [EOS] tokens to indicate sequence boundaries. The resulting token sequence $\{f_{sos}^{t},f_{1}^{t},\dots,f_{eos}^{t}\}$ is processed by the transformer encoder, which models dependencies among tokens via masked self-attention. Finally, the [EOS] token embedding from the top layer, $f_{eos}^{t}$, is linearly mapped into the joint image-text representation space, generating a compressed global textual representation.
@@ 3.2. Adaptive Similarity Distribution Matching
Adopted from IRRA (jiang2023cross), we introduce a novel Adaptive Similarity Distribution Matching (A-SDM) module, which aims to adaptively pull the unmatched positive image-text pairs into a shared representation space, further enhancing the cross-modal global matching capability of the model.
Let the mini-batch contain $B$ image-text pairs, we pair each text embedding $g_{i}^{t}$ with its global image embedding $g_{j}^{v}$ to form the set $\{(g_{i}^{t},g_{j}^{v}),y_{i,j}\}_{j=1}^{B}$, where $y_{i,j}$ serves as the matching indicator. Specifically, $y_{i,j}=1$ denotes a matched pair, while $y_{i,j}=0$ denotes an mismatched pair. Let $cos(\mathbf{a},\mathbf{c})=\mathbf{a}^{\top}\mathbf{c}/\|\mathbf{a}\|\|\mathbf{c}\|$ denotes the similarity of $\mathbf{a}$ and $\mathbf{c}$. Subsequently, like SDM (jiang2023cross), the similarity matrix of image-text pairs is obtained through the following softmax function:
(1) $p_{i,j}=\frac{exp(cos(g_{i}^{t},g_{j}^{v})/\tau_{1})}{\sum_{k=1}^{B}exp(cos(g_{i}^{t},g_{k}^{v})/\tau_{1})},$
where $\tau_{1}$ acts as a temperature term that modulates the spread of the resulting distribution. The probability $p_{i,j}$ quantifies how much the similarity between the text embedding $g_{i}^{t}$ and the image embedding $g_{j}^{v}$ contributes relative to the sum of all similarities between $g_{i}^{t}$ and every image embedding in the mini-batch.
Let the ithi-th text $T_{i}$ from the batch be designated as the query text and $I_{i}$ be the corresponding image for $T_{i}$ at rank-$k$, where $k>1$. Different from IRRA (jiang2023cross), we propose to derive an adaptive weighting factor by assessing the similarity between the query text $T_{i}$ and all image representations:
@@
where $\alpha$ is a weight factor reflecting the contribution of unmatched image-text pairs to the cross-modal global matching ability of the model. Here, $\max_{k}p_{i,k}$ indicates the top similarity value between the text $T_{i}$ and every image within the mini-batch, while $p_{i,i}$ refers to the similarity associated with its corresponding positive image. The constant term “+1+1” ensures that when $T_{i}$ and its corresponding image $I_{i}$ are correctly matched, the weight $w_{i}^{t2i}$ defaults to 1. In this case, the A-SDM loss reduces to the SDM loss (jiang2023cross), preventing overemphasis on correctly matched pairs while allowing the model to focus adaptively on harder and misaligned pairs. Conversely, $w_{i}^{t2i}>1$ indicates that $T_{i}$ and $I_{i}$ are unmatched, increasing their contribution to the loss to enhance global cross-modal alignment. The A-SDM loss for mapping text to image within a mini-batch is subsequently formulated as:
@@
where $\epsilon$ is a tiny offset added to safeguard the computation from unstable values, and $q_{i,j}=y_{i,j}/\sum_{k=1}^{B}$ denotes the ground-truth matching probability.
In a complementary manner, the A-SDM loss for the image-to-text branch $\mathcal{L}_{i2t}$ is derived by swapping the roles of the text and image features. The bi-directional A-SDM loss is formulated as: